# Rethinking Data Augmentation for Robust LiDAR Semantic Segmentation in Adverse Weather

### v2

@@ Abstract
Existing LiDAR semantic segmentation methods often struggle with performance declines in adverse weather conditions. Previous research has addressed this issue by simulating adverse weather or employing universal data augmentation during training. However, these methods lack a detailed analysis and understanding of how adverse weather negatively affects LiDAR semantic segmentation performance. Motivated by this issue, we identified key factors of adverse weather and conducted a toy experiment to pinpoint the main causes of performance degradation: (1) Geometric perturbation due to refraction caused by fog or droplets in the air and (2) Point drop due to energy absorption and occlusions. Based on these findings, we propose new strategic data augmentation techniques. First, we introduced a Selective Jittering (SJ) that jitters points in the random range of depth (or angle) to mimic geometric perturbation. Additionally, we developed a Learnable Point Drop (LPD) to learn vulnerable erase patterns with a Deep Q-Learning Network to approximate the point drop phenomenon from adverse weather conditions. Without precise weather simulation, these techniques strengthen the LiDAR semantic segmentation model by exposing it to vulnerable conditions identified by our data-centric analysis. Experimental results confirmed the suitability of the proposed data augmentation methods for enhancing robustness against adverse weather conditions. Our method attains a remarkable 39.5 mIoU on the SemanticKITTI-to-SemanticSTF benchmark, surpassing the previous state-of-the-art by over 5.4%p, tripling the improvement over the baseline compared to previous methods achieved. Our code will be released at https://github.com/engineerJPark/LiDAR-DataAug4Weather.
@@
Finally, we apply the proposed augmentations when training the LiDAR segmentation model with a clean source dataset. Experimental results demonstrate that our method achieves new state-of-the-art performance on SemanticSTF, a real adverse weather dataset. Remarkably, we achieved an impressive +8.1 mIoU gain over the baseline model, MinkowskiNet. This is a notable improvement, nearly tripling the +2.5 mIoU improvement from previous work [33]. Additionally, we observed performance gains of +7.8 and +10.3 mIoU for CENet and SPVCNN, respectively, in SemanticSTF. Our method also shows performance gains in SemanticKITTI-C, with +3.9, +0.4, and +5.6 mIoU for CENet, SPVCNN, and MinkowskiNet, respectively. Furthermore, our method outperforms recent domain generalization methods, achieving an increase of +4.9 mIoU.
@@ 3.2 Toy Experiment
Based on previous research findings mentioned in Section 3, the types of distortions in LiDAR point clouds caused by adverse weather conditions converge into a set of common distortions. Therefore, from a data-centric perspective, the issues that we need to consider can be summarized as follows: (D1) Point drop, (D2) Occlusions, (D3) Geometric perturbation, and (D4) Intensity distortion. Here, we aim to identify which distortion types negatively impact performance. To achieve this, we have generated four distortion types of toy synthetic data through the SemanticKITTI validation set.
@@
(D2) Occluding Points: The method assumes that occlusions occur predominantly in front of objects, mainly due to distortion from fog, snow, and rain. We synthesized this data by randomly selecting points and altering their depth to one-tenth of the original depth. The selecting ratios were also determined to be 0.5 and 0.9.
@@ 6 Conclusion
Acknowledgements. This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.2021-0-02068, Artificial Intelligence Innovation Hub and RS-2019-II190075, Artificial Intelligence Graduate School Program(KAIST)), IITP grant funded by the Korea government(MSIT) and KEIT grant funded by the Korea government(MOTIE) (No.2022-0-00680, No.2022-0-01045), Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the MSIP (NRF-2022R1A2C3011154, RS-2023-00219019, RS-2023-00240135) and Samsung Electronics Co., Ltd (IO230508-06190-01).

### v4

# supplementary.pdf

@@ Abstract
Existing LiDAR semantic segmentation methods often struggle with performance declines in adverse weather conditions. Previous work has addressed this issue by simulating adverse weather or employing universal data augmentation during training. However, these methods lack a detailed analysis and understanding of how adverse weather negatively affects LiDAR semantic segmentation performance. Motivated by this issue, we identified key factors of adverse weather and conducted a toy experiment to pinpoint the main causes of performance degradation: (1) Geometric perturbation due to refraction caused by fog or droplets in the air and (2) Point drop due to energy absorption and occlusions. Based on these findings, we propose new strategic data augmentation techniques. First, we introduced a Selective Jittering (SJ) that jitters points in the random range of depth (or angle) to mimic geometric perturbation. Additionally, we developed a Learnable Point Drop (LPD) to learn vulnerable erase patterns with a Deep Q-Learning Network to approximate the point drop phenomenon from adverse weather conditions. Without precise weather simulation, these techniques strengthen the LiDAR semantic segmentation model by exposing it to vulnerable conditions identified by our data-centric analysis. Experimental results confirmed the suitability of the proposed data augmentation methods for enhancing robustness against adverse weather conditions. Our method achieves a notable 39.5 mIoU on the SemanticKITTI-to-SemanticSTF benchmark, improving the baseline by 8.1%p and establishing a new state-of-the-art. Our code will be released at https://github.com/engineerJPark/LiDARWeather.
@@
Finally, we apply the proposed augmentations when training the LiDAR segmentation model with a clean source dataset. Experimental results demonstrate that our method achieves new state-of-the-art performance on SemanticSTF, a real adverse weather dataset. Specifically, we achieved an impressive +8.1%p mIoU gain over the baseline MinkowskiNet, nearly tripling the +2.5%p mIoU improvement from previous work [33]. Our method achieved remarkable performance gains, ranging from +0.4 to +10.3%p mIoU, across various model architectures (Minkowski, CENet, SPVCNN) and corruption dataset (SemanticSTF and SemanticKITTI-C). These results demonstrate the generalizability of the proposed augmentation methods.
@@ 3.2 Toy Experiment
Based on previous research findings mentioned in Section 3.1, the types of distortions in LiDAR point clouds caused by adverse weather conditions converge into a set of common distortions. Therefore, from a data-centric perspective, the issues that we need to consider can be summarized as follows: (D1) Point Drop, (D2) Occlusions, (D3) Geometric Perturbation, and (D4) Intensity Distortion. Here, we aim to identify which distortion types negatively impact performance. To achieve this, we have generated four distortion types of toy synthetic data through the SemanticKITTI validation set.
@@
(D2) Occlusions: The method assumes that occlusions occur predominantly in front of objects, mainly due to distortion from fog, snow, and rain. We synthesized this data by randomly selecting points and altering their depth to one-tenth of the original depth. The selecting ratios were also determined to be 0.5 and 0.9.
@@ 6 Conclusion
Acknowledgements. This work was supported by IITP grant funded by MSIT (No.2021-0-02068, Artificial Intelligence Innovation Hub and RS-2019-II190075, Artificial Intelligence Graduate School Program(KAIST)) and KEIT grant funded by MOTIE (No.2022-0-00680, No.2022-0-01045), NRF funded by the MSIP (NRF-2022R1A2C3011154, RS-2023-00219019, RS-2023-00240135) and Samsung Electronics Co., Ltd (IO230508-06190-01).